OpenStudy (anonymous):

the perimeter of a rectangle is 54cm the area of the same rectangle is 176 cm^2. what are the dimensions of the rectangle

OpenStudy (anonymous):

a) 11 cm by 16 cm b) 8cm by 22 cm c) 5.5 cm by 32 cm d) 4 cm by 44 can someone teach me how to do it?

OpenStudy (anonymous):

Ok so first of all , lets assume the dimensions are X and Y , then you have to equations P=54 >>> 2(X+Y)=54 >>> X+Y=27 ...(1) A=176 >>> XY=176 ...(2) now in equation 1 you can solve for Y and you get Y=27-X and substitute it in equation 2 and solve for X at the end u will get the value of X to be 16 or 11 , when X=16 then Y=11 , OR when X=11 then Y=16

OpenStudy (anonymous):

so its 11 cm by 16 cm?